# Genetic-guided GFlowNets for Sample Efficient Molecular Optimization

### v2

@@ 1 Introduction
To address this, we propose a novel molecular optimization algorithm that integrates domain-specialized genetic algorithms into GFlowNets, an off-policy training method that trains policy to sample proportional to their rewards. We first generate diverse candidate samples using the current policy and refine these candidates into higher-reward samples using GAs (see Fig. 1). Subsequently, we fine-tune the policy with a GFlowNet using the collected samples. Unlike the MLE training, which can be stuck in local modes when the reward landscape is peaky, GFlowNet trains the policy to sample molecules proportional to rewards. To enhance sample efficiency, we perform unsupervised pretraining on chemical datasets and regularize the GFlowNets policy with KL-divergence to align generative probability with the dataset distribution, focusing on the compact valid space.
@@
  • GA increases the exploitation power of GFlowNets. The proposed algorithm incorporates an effective off-policy exploration into GFlowNets based on domain-specialized GA. This approach aligns with recent studies that utilize off-policy explorations to guide toward the high-reward regions [18]. Our key contribution lies in explicitly leveraging domain knowledge about chemical structures and effectively distilling it into GFlowNets, which enable exceptional performance in real-world tasks beyond small-scale molecular generations. This contribution is crucial for the field of GFlowNets, which have struggled with sample-efficient molecular optimization tasks, even when using proxy reward models for active learning [10, 8]; see experiments in Section 4.2.
  • GFlowNets increases population diversity of GA. The proposed algorithm generates diversified samples using GFlowNets, enabling GA to effectively improve samples; our method can be regarded as a new genetic algorithm with deep generative policy-based population resetting. Our GFlowNet policy, parameterized over the whole space, periodically resets the population by diversely sampling individuals proportional to rewards, mitigating premature convergence of GA [19, 20]. Experiments show that ours outperforms recent GAs in PMO benchmark; see Section 5.1.
@@ 2.2 Generative flow networks
In GFlowNet training, employing exploratory behavior policies or replay training is allowed since GFlowNet can be trained in an off-policy manner, which is a key advantage [23, 24, 28, 18, 29].
@@ 3.1 Factorized string-based generative policy and unsupervised pretraining
Building on insights from previous works, we employ a string-based representation strategy, simplifying the molecular generation process by reducing it to a one-directional sequence generation. We adopt a sequence generative policy using a string-based assembly strategy, especially the simplified molecular-input line-entry system (SMILES) [30]. Motivated by REINVENT [3], we parameterize the policy using a recurrent neural network architecture [31]. Then, the probability $\pi_{\theta}(\bm{x})$ of generating a molecule, can be factorized to $\prod_{t=1}^{n}\pi_{\theta}(x_{t}|x_{1},\ldots,x_{t-1})$, where $x_{1},\ldots,x_{n}$ are characters of SMILES representation of $\bm{x}$.
@@ 3.2.2 Updating the generative policy with GFlowNets training
The generative policy is trained using the trajectory balance loss in Eq. 1. Note that we set $P_{B}$ to 1 for simplicity since SMILES generations are conducted in one direction. To ensure that the policy does not deviate excessively from the pretrained policy during training, we introduce a Kullback-Leibler divergence penalty inspired by the works in language model fine-tuning [32, 33]. Thus, our model is updated to minimize the following loss function:
@@ 4.2 GFlowNets for molecular optimization
Generative Flow Networks (GFlowNets or GFN) [23, 6] have drawn significant attention in scientific discovery [38], particularly in molecular optimization and biological sequence design [6, 10, 39, 18, 40, 41, 11]. GFlowNets, which are off-policy variational inference methods [42], are closely related to value-based reinforcement learning within soft Markov Decision Processes (soft MDPs) [43, 44], focusing on learning maximum entropy agents [45, 46]. This allows GFlowNets to generate diverse candidates for chemical and biological structures. Advances in GFlowNets have included new objective functions [6, 24, 25], improved credit assignments [26, 27], and enhanced off-policy exploration strategies [47, 18, 48, 49]. Our approach, an improved off-policy exploration technique for GFlowNets, enhances sample efficiency and scalability for moderate-scale chemical discovery. We present extensive experiments targeting the discovery of larger-scale molecules, with SMILES strings approximately 100 characters long, surpassing the scope of smaller molecule generation tasks commonly addressed in GFlowNets literature [6, 39].
@@ 5 Experiments
5.1 Sample efficient molecular optimization
5.1.1 Main results in the official benchmark of PMO
Table 1: Mean and standard deviation of AUC top-10 ($\uparrow$) from five independent runs. We use oracle ID (in lexicographical order) instead of a name for better readability, and the best mean scores are denoted in bold for each task. The results of further baselines are provided in Section F.2.
@@ 5.1.1 Main results in the official benchmark of PMO
Figure 2: The optimization curve of the average scores of Top-10 over the score function calls. All optimization curves for 23 oracles are provided in Section F.2.
@@
Moreover, we adopt additional methods, Mol GA [16] and GEGL [17]. Mol GA is an advanced version of Graph GA and outperforms other baselines in the PMO benchmark. On the other hand, GEGL is an ablated version of our approach that utilizes imitation learning with a reward-priority queue instead of GFlowNet training with rank-based sampling. For both, we adopt the original implementations [^1] https://github.com/AustinT/mol_ga [^2] https://github.com/sungsoo-ahn/genetic-expert-guided-learning with hyperparameters searches following the guidelines; see Appendix B.
The main results in Table 1 report the AUC score of Top-10 candidates with independent five runs with different seeds. In addition, Fig. 2 visually presents the Top-10 average score across the computational budget, i.e., the number of oracle calls, providing a concise overview of the results. Due to the lack of space, the best five results are provided; please check Section F.2 for the rest of the results. As shown in Table 1, Genetic GFN outperforms the other baselines with a total of 16.213 and attains the highest AUC Top-10 values in 14 out of 23 score functions. The results of diversity and SA score for each oracle are presented in Section F.4.
@@ 5.1.2 Controllability of the scores-diversity trade-off and ablation studies
In the benchmark, we found a pronounced trade-off between attaining high evaluation scores within a limited budget and generating diverse molecular candidates. This section demonstrates the controllability of the score-diversity trade-off through adjustments in the inverse temperature $\beta$. Decreasing the inverse temperature gives more diverse candidates. The results in Fig. 3 demonstrate adjustments of $\beta$ can control the trade-off between score and diversity, achieving Pareto-frontier to other baselines in the benchmark. Notably, Genetic GFN with $\beta=30$ achieves a higher AUC Top-10 with a greater diversity compared to the SOTA GA method (Mol GA: 15.686 with a diversity of 0.465) and RL method (REINVENT: 15.185 with a diversity of 0.468). Similarly, the weight-shifting factor $k$ in rank-based sampling can control the trade-off; see Section F.1.
@@
Ablation studies. The ablation studies investigate the essential components of our framework: the genetic search (GS) and the KL-divergence penalty. To assess the effectiveness of the genetic search, we also compare its performance against the exploration strategy used in previous GFlowNet studies [6, 10]. It samples actions from a GFlowNet sampler mixed with a uniform distribution, similar to $\epsilon$-greedy in RL. The results, shown in Table 2, reveal that the removal of either component results in a decline in performance, underscoring the importance of employing a suitable exploration strategy. Detailed results, including statistical analysis, are provided in Section F.3.
@@ 5.1.4 Sample efficient multi-objective molecular optimization
Table 4: Average and standard deviation of hypervolumes ($\uparrow$) for each task. The baseline results are directly from the HN-GFN paper. The bold text indicates the best value.
@@ 5.1.5 Further analysis
Active learning with Genetic GFN. Similar to GFlowNet-AL, ours can work as a generative model in multi-round active learning. We compare Genetic GFN-AL with other model-based and active learning methods; please refer to Appendix C.
Genetic GFN with SELFIES representation. Genetic GFN with SELFIES generation achieves the improved sample efficiency to other SELFIES-based methods; see Section F.5.
Sensitivity analysis. We provide the experimental results by varying the hyperparameters, such as the offspring size and the number of training loops; see Section F.6.
@@ 5.2 Designing inhibitors against SARS-CoV-2 targets
The experiments are conducted with up to 1000 update steps with 128 batch size [52]. As shown in Table 5, ours achieves the highest Top-100 average scores only with 100 steps, which is 10 times fewer than others. Note that the score is recalculated based on the normalized score function in Eq. 5 using average values in the MolRL-MGPT paper [52]; the full results and a more detailed experimental setup are provided in Appendix E. We also report the best candidates of 100 steps in Fig. 4 and in Fig. 5. The final molecules correspond to the Top-1 score molecules from 3 independent runs.
@@ 6 Discussion
This paper introduces a Genetic-guided GFlowNet (Genetic GFN), which integrates a domain-specific genetic algorithm to guide the GFlowNet policy toward higher-reward samples. The method employs off-policy training with a rank-based reweighted buffer, enhancing the policy as a powerful amortized inference sampler for chemical discovery. Extensive experiments demonstrate that Genetic GFN effectively generates desirable molecules within the high-dimensional chemical space, including long chemical structure sequences (e.g., $\geq$ 100). On the other hand, our approach can be considered as a novel population reinitialization strategy for genetic algorithms using GFlowNets, which samples diverse objects proportional to rewards.
Limitations and future works. Our method assumes the existence of effective genetic algorithms, which is valid for the molecular design domain. However, designing domain-specific operators for genetic algorithms can be challenging in other fields. One possible future work is to enhance genetic algorithms using the neural policy similar to recent studies [54, 55]. Another direction is to extend our approach to other domains, such as combinatorial optimization. For example, our algorithm exhibits similarities with recent developments in neural combinatorial optimization [56, 57, 58, 59, 60], where the neural policy samples solution trajectories in a constructive process. As pointed out in the recent work [61], CO problems in real-world applications often involve expensive reward evaluations. In this context, we could utilize a powerful GA, such as hybrid genetic search (HGS) [62], to further enhance the sample efficiency of these neural combinatorial optimization solvers.
@@ A.1 Network architecture and pretraining
Pre-training.
According to the PMO benchmark guidelines [8], the pre-training is conducted on ZINC 250K. Since the network architecture is the same as REINVENT, we adopt the provided pretrained model for REINVENT in the PMO benchmark. This allows the direct comparison of fine-tuning approaches.
@@ A.2 Hyperparameters
In constrast, we have searched several hyperparameters, offspring size, the number of GA generations, and KL-divergence coefficient $\alpha$. We provide the sensitivity analysis for the offspring size and the number of GA generations in Section F.6. Furthermore, we use the inverse temperature $\beta=10$ and the weight-shifting factor $k=0.01$, but they can be differently used to control score-diversity trade-off, as explained in Section 5.1.2.
@@ A.3 Computing resource
Appendix B Hyperparameter setup for baseline methods
@@
Appendix C Comparison Genetic GFN-AL with active learning and model-based algorithms
C.1 Implementation of Genetic GFN-AL

### v4

@@ 1 Introduction
To address this, we propose a novel molecular optimization algorithm that integrates domain-specialized genetic algorithms into GFlowNets, an off-policy training method that trains policy to sample proportional to their rewards. As illustrated in Fig. 1, we first generate diverse candidates using the current policy and refine the candidates into higher-reward samples using GAs. Subsequently, we fine-tune the policy with a GFlowNet using the collected samples. Unlike the MLE training, which can be stuck in local modes when the reward landscape is peaky, GFlowNet trains the policy to sample molecules proportional to rewards. To enhance sample efficiency, we perform unsupervised pretraining on chemical datasets and regularize the GFlowNets policy with KL-divergence to align generative probability with the dataset distribution, focusing on the compact valid space.
@@
  • GA increases the exploitation power of GFlowNets. The proposed algorithm incorporates an effective off-policy exploration into GFlowNets based on domain-specialized GA. This approach aligns with recent studies that utilize off-policy explorations to guide toward the high-reward regions [18]. Our key contribution lies in explicitly leveraging domain knowledge about chemical structures and effectively distilling it into GFlowNets, which enable exceptional performance in real-world tasks beyond small-scale molecular generations. This contribution is crucial for the field of GFlowNets, which have struggled with sample-efficient molecular optimization tasks, even when using proxy reward models for active learning [10, 8]; see Section 4.2.
  • GFlowNets increases population diversity of GA. The proposed algorithm generates diversified samples using GFlowNets, enabling GA to effectively improve samples; our method can be regarded as a new genetic algorithm with deep generative policy-based population resetting. Our GFlowNet policy, parameterized over the whole space, periodically resets the population by diversely sampling individuals proportional to rewards, mitigating premature convergence of GA [19, 20]. Experiments show that ours outperforms recent GAs in the PMO benchmark; see Section 5.1.
@@ 2.2 Generative flow networks
In GFlowNet training, employing exploratory behavior policies or replay training is allowed since GFlowNet can be trained in an off-policy manner, which is a key advantage [23, 24, 28, 18].
@@ 3.1 Factorized string-based generative policy and unsupervised pretraining
Building on insights from previous works [3, 17], we employ a string-based representation strategy, simplifying the molecular generation process by reducing it to a one-directional sequence generation. We adopt a sequence generative policy using a string-based assembly strategy, especially the simplified molecular-input line-entry system (SMILES) [29]. Motivated by REINVENT [3], we parameterize the policy using a recurrent neural network architecture [30]. Then, the probability $\pi_{\theta}(\bm{x})$ of generating a molecule, can be factorized to $\prod_{t=1}^{n}\pi_{\theta}(x_{t}|x_{1},\ldots,x_{t-1})$, where $x_{1},\ldots,x_{n}$ are characters of SMILES representation of $\bm{x}$.
@@ 3.2.2 Updating the generative policy with GFlowNets training
The generative policy is trained using the trajectory balance loss in Eq. 1. Note that we set $P_{B}$ to 1 for simplicity since SMILES generations are conducted in one direction. To ensure that the policy does not deviate excessively from the pretrained policy during training, we introduce a Kullback-Leibler (KL) divergence penalty inspired by the works in language model fine-tuning [31, 32]. Thus, our model is updated to minimize the following loss function:
@@ 4.2 GFlowNets for molecular optimization
Generative Flow Networks (GFlowNets or GFN) [23, 6] have drawn significant attention in scientific discovery [37], particularly in molecular optimization and biological sequence design [6, 10, 38, 18, 39, 40, 11, 41]. GFlowNets, which are off-policy variational inference methods [42], are closely related to value-based reinforcement learning within soft Markov Decision Processes (soft MDPs) [43, 44], focusing on learning maximum entropy agents [45, 46]. This allows GFlowNets to generate diverse candidates for chemical and biological structures. Advances in GFlowNets have included new objective functions [6, 24, 25], improved credit assignments [26, 27], and enhanced off-policy exploration strategies [18, 47, 48, 49]. Our approach, an improved off-policy exploration technique for GFlowNets, enhances sample efficiency and scalability for moderate-scale chemical discovery. We present extensive experiments targeting the discovery of larger-scale molecules, with SMILES strings approximately 100 characters long, surpassing the scope of smaller molecule generation tasks commonly addressed in GFlowNets literature [6, 38].
@@ 5 Experiments
Table 1: Mean and standard deviation of AUC top-10 ($\uparrow$) from five independent runs. We use oracle ID (in lexicographical order) instead of a name for better readability, and the best mean scores are denoted in bold for each task. The results of further baselines are provided in Section G.2.
@@
Figure 2: The optimization curve of the average scores of Top-10 over the score function calls. All optimization curves for 23 oracles are provided in Section G.2.
5.1 Sample efficient molecular optimization
5.1.1 Main results in the official benchmark of PMO
@@
Moreover, we adopt additional methods, Mol GA [16] and GEGL [17]. Mol GA is an advanced version of Graph GA and outperforms other baselines in the PMO benchmark. On the other hand, GEGL is an ablated version of our approach that utilizes imitation learning with a reward-priority queue instead of GFlowNet training with rank-based sampling. For both, we adopt the original implementations [^1] https://github.com/AustinT/mol_ga [^2] https://github.com/sungsoo-ahn/genetic-expert-guided-learning with hyperparameters searches following the guidelines; see Appendix C.
The main results in Table 1 report the AUC score of Top-10 candidates with independent five runs with different seeds. In addition, Fig. 2 visually presents the Top-10 average score across the computational budget, i.e., the number of oracle calls, providing a concise overview of the results. Due to the lack of space, the best five results are provided; please check Section G.2 for the rest of the results. As shown in Table 1, Genetic GFN outperforms the other baselines with a total of 16.213 and attains the highest AUC Top-10 values in 14 out of 23 score functions. The results of diversity and SA score for each oracle are presented in Section G.5.
@@ 5.1.2 Controllability of the scores-diversity trade-off and ablation studies
In the benchmark, we found a pronounced trade-off between attaining high evaluation scores within a limited budget and generating diverse molecular candidates. This section demonstrates the controllability of the score-diversity trade-off through adjustments in the inverse temperature $\beta$. Decreasing the inverse temperature gives more diverse candidates. The results in Fig. 3 demonstrate adjustments of $\beta$ can control the trade-off between score and diversity, achieving Pareto-frontier to other baselines in the benchmark. Notably, Genetic GFN with $\beta=30$ achieves a higher AUC Top-10 with a greater diversity compared to the SOTA GA method (Mol GA: 15.686 with a diversity of 0.465) and RL method (REINVENT: 15.185 with a diversity of 0.468). Similarly, the weight-shifting factor $k$ in rank-based sampling can control the trade-off; see Section G.1.
@@
Ablation studies. The ablation studies investigate the essential components of our framework: the genetic search (GS) and the KL-divergence penalty. To assess the effectiveness of the genetic search, we also compare its performance against the exploration strategy used in previous GFlowNet studies [6, 10]. It samples actions from a GFlowNet sampler mixed with a uniform distribution, similar to $\epsilon$-greedy in RL. The results, shown in Table 2, reveal that the removal of either component results in a decline in performance, underscoring the importance of employing a suitable exploration strategy. Detailed results, including statistical analysis, are provided in Section G.4.
@@ 5.1.4 Sample efficient multi-objective molecular optimization
Table 4: Average and standard deviation of hypervolumes ($\uparrow$) for each task. The baseline results are directly from the HN-GFN paper [11]. The bold text indicates the best value.
@@ 5.1.5 Further analysis
Active learning with Genetic GFN. Similar to GFlowNet-AL, ours can work as a generative model in multi-round active learning. We compare Genetic GFN-AL with other model-based and active learning methods; please refer to Appendix D.
Genetic GFN with SELFIES representation. Genetic GFN with SELFIES generation achieves the improved sample efficiency to other SELFIES-based methods; see Section G.6.
Sensitivity analysis. We provide the experimental results by varying the hyperparameters, such as the offspring size and the number of training loops; please refer to Section G.8.
@@ 5.2 Designing inhibitors against SARS-CoV-2 targets
The experiments are conducted with up to 1000 update steps with 128 batch size [52]. As shown in Table 5, ours achieves the highest Top-100 average scores only with 100 steps, which is 10 times fewer than others. Note that the score is recalculated based on the normalized score function in Eq. 5 using average values in the MolRL-MGPT paper [52]; the full results and a more detailed experimental setup are provided in Appendix F. We also report the best candidates of 100 steps in Fig. 4 and in Fig. 5. The final molecules correspond to the Top-1 score molecules from 3 independent runs.
@@ 6 Discussion
This paper introduces a Genetic-guided GFlowNet (Genetic GFN), which integrates a domain-specific genetic algorithm to guide the GFlowNet policy toward higher-reward samples. The method employs off-policy training with a rank-based reweighted buffer, enhancing the policy as a powerful amortized inference sampler for chemical discovery. Extensive experiments demonstrate that Genetic GFN effectively generates desirable molecules within the high-dimensional chemical space, including long chemical structure sequences (e.g., $\geq$ 100). On the other hand, our approach can be considered as a novel population reinitialization strategy for genetic algorithms using GFlowNets, which sample diverse objects proportional to rewards.
Limitations and future works. Our method assumes the existence of effective genetic algorithms, which is valid for the molecular design domain. However, designing domain-specific operators for genetic algorithms can be challenging in other fields. One possible future work is to enhance genetic algorithms using the neural policy similar to recent studies [54, 55]. Another direction is to extend our approach to other domains, such as combinatorial optimization. For instance, we could utilize a powerful GA, hybrid genetic search [56], to design a GFlowNet-based solver for routing problems.
@@
Acknowledgments and Disclosure of Funding
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00410082). The authors are grateful to Austin Tripp and Xiuyuan Hu for their help with baselines.
@@ Appendix A Implementation details of Genetic GFN
Figure 6: Pretraining and Genetic GFN fine-tuning framework
@@ A.1 Network architecture and pretraining
Pretraining.
According to the PMO benchmark guidelines [8], the pre-training is conducted on ZINC 250K. The overall framework is illustrated in Fig. 6. Since the network architecture is the same as REINVENT, we adopt the provided pretrained model for REINVENT in the PMO benchmark. This allows the direct comparison of fine-tuning approaches.
@@ A.2 Hyperparameters
In contrast, we have searched several hyperparameters, offspring size, the number of GA generations, and KL-divergence coefficient $\alpha$. We provide the sensitivity analysis for the offspring size and the number of GA generations in Section G.8. Furthermore, we use the inverse temperature $\beta=10$ and the weight-shifting factor $k=0.01$, but they can be differently used to control score-diversity trade-off, as explained in Section 5.1.2.
@@ A.3 Computing resource
Appendix B Genetic operations
This section details each operation in our genetic search; see Fig. 7 for illustration. Note that we adopt Graph GA of [13], which has demonstrated its powerful performances and has been adopted by GA-related works like Mol GA [16] and GEGL [17].
(a) Crossover
(b) Mutations
Figure 7: Examples of Graph GA operations. These operations are conducted according to predefined SMARTS patterns to ensure molecule validity, such as adherence to valence rules.
B.1 Crossover
A crossover operation is conducted to generate a new candidate (called offspring) by exchanging the genetic information of a pair of selected individuals (parents). This process mimics the crossover of genetic material in biological reproduction. In the context of molecular optimization with graphs, the crossover operation is conducted in two types: ‘ring crossover’ and ‘non-ring crossover with a 50% probability.
These two main crossover operations perform crossover between two parent molecules by cutting and recombining ring substructures. Ring crossover performs a ring cut specifically designed to target ring structures within the molecule. The ring-cut operation cuts the molecule along two different ring patterns, selected randomly. One of the ring patterns checks for a specific arrangement of four consecutive ring atoms, and the other pattern checks for a ring atom connected to two other ring atoms with a single bond. If a suitable ring pattern is found, it cuts the molecule along that pattern, resulting in two fragments. On the other hand, non-ring crossover cuts a single bond, meaning it is not part of a cyclic (ring) structure within the molecule. The obtained fragments from both parents are recombined to create new molecules by applying predefined reaction SMARTS patterns. These operations are repeated for validity to ensure that the resulting molecules meet structural and size constraints.
B.2 Mutation
The mutation is a random change that is introduced to the genetic information of some individuals. This step adds diversity to the population and helps explore new regions of the solution space. In this work, we employ seven different mutation processes and randomly select one of these mutations to modify the offspring molecules slightly. The operations consist of atom and bond deletions, appending new atoms, inserting atoms between existing ones, changing bond orders, deleting cyclic bonds, adding cyclic rings, and altering atom types.
  1. 1.
    Deletion of atom: it selects one of five deletion SMARTS patterns, each representing the removal of a specific number of atoms or bonds. These patterns include the removal of a single atom, a single bond, a bond with two attached atoms, and bonds with multiple attached atoms. The selected pattern is applied to the molecule, deleting the specified atom(s) or bond(s).
  2. 2.
    Appending atom: it introduces a new atom to the molecule. The type of atom (e.g., C, N, O) and the type of bond (single, double, or triple) are chosen based on predefined probabilities. The function then generates a reaction SMARTS pattern to append the selected atom to the molecule, forming a new bond.
  3. 3.
    Inserting atom: it inserts a new atom between two existing atoms in the molecule. Similar to the appending atom, it selects the type of atom and bond based on predefined probabilities and generates a reaction SMARTS pattern to insert the atom.
  4. 4.
    Changing bond order: it randomly selects one of four SMARTS patterns, each representing a change in the bond order between two atoms. These patterns include changing a single bond to a double bond, a double bond to a triple bond, and vice versa.
  5. 5.
    Deletion of cyclic bond: it deletes a bond that is a part of a cyclic structure within the molecule. The SMARTS pattern represents the breaking of a cyclic bond while retaining the atoms connected by the bond.
  6. 6.
    Adding ring: it introduces a new cyclic ring into the molecule by selecting one of four SMARTS patterns, each representing the formation of a specific ring type. These patterns create different types of cyclic structures within the molecule.
  7. 7.
    Changing atom: it randomly selects two types of atoms from a predefined list and generates a SMARTS pattern to change one atom into another. This operation modifies the atom type within the molecule.
Appendix C Hyperparameter setup for baseline methods
@@
Appendix D Comparison Genetic GFN-AL with active learning and model-based algorithms
D.1 Implementation of Genetic GFN-AL